OpenStudy (anonymous):

If f(x) = −4x + 1, find f(−5)."

OpenStudy (anonymous):

to find f(-5) simply plug -5 in where-ever there is an "X" f(-5)=-4(-5)+1 =20+1 =21 winner winner chicken dinner! ;D

OpenStudy (anonymous):

hey is this a subset N: {1,2,3,4,5} {1,2,3} {2} {1,2,3,4,5} {}